OpenStudy (anonymous):

Converting a Riemann sum to a definite integral. So, I'm supposed to find \[\lim_{n \to \infty} \sum_{i=1}^n \frac{sin(i \pi /n)}{n}\] by evaluating an appropriate definite integral over the interval [0, 1]. Here's my work so far: \[\lim_{n \to \infty} \sum_{i=1}^n \frac{sin(i \pi /n)}{n} = \lim_{n \to \infty} \sum_{i=1}^n \left[ \sin \left(\frac{i \pi}{n} \right) \cdot \frac{1}{n}\right]\] \[= \int_0^1\! sin(x) dx\] \[= 1 - cos(1)\] No clue if I'm on the right track here, any help is appreciated.

ganeshie8 (ganeshie8):

integrand should be \(\large \sin(\color{red}{\pi} x)\) right ?

OpenStudy (anonymous):

Is that what I missed? I was wondering where to put that \( \pi\) symbol.

ganeshie8 (ganeshie8):

Yes only that, everything else looks perfect!

OpenStudy (anonymous):

Sweet! Thank you, Ganeshie8 :)

ganeshie8 (ganeshie8):

you're welcome :)

ganeshie8 (ganeshie8):

wolfram is your best friend in double checking the final answer http://www.wolframalpha.com/input/?i=lim+n%5Cto+%5Cinfty+%5Csum%5Climits_%7Bi%3D1%7D%5E%7Bn%7Dsin%28ipi%2Fn%29%2Fn+

OpenStudy (anonymous):

Ah sweet, I tried to use wolframalpha to check but I couldn't figure out how to input the expression; I didn't know you could basically input the LaTeX code.

ganeshie8 (ganeshie8):

yes copy pasting the latex code into wolfram works almost all the time !

OpenStudy (anonymous):

Sweet! that'll help with checking a lot of my answers, thank you.